So, we started learning about systems of equations, and I'm a bit tripped up.
{4X+2Y=6
{-3X-7Y=1
Using substitution how would I solve this system?
Another one that has me confused is this:
{X+Y=13
{X-Y=2
I have tried it a couple times before, but I'm quite sure they aren't correct.

Answer by mananth(16946)

4x+2y=6
/2
we get 2x+y=3
2 x + y = 3 .............1
y = 3 + -2 x

y = 3 -2 x
-3 x -7 y = 1 .....................2
Substitute y in (2)
-3 x -7 ( 3 -2 x ) = 1
-3 x -21 + 14 x = 1
-3 x 14 x = 22
11 x = 22
/ 11
x= 2
Plug the value of x in (1)
2 x + y = 3
2 * 2 + 1 y = 3
4 + 1 y = 3
y = -1
/ 1
y= -1
--------------
x+y=13
x-y=2
in the second equation
y=x-2substitute y in (1)
x+(x-2)=13
2x-2=13
2x=15
x=15/2
Plug x in (2)
x-y=2
(15/2)-y=2
y= 15/2 -2
y=11/2